# QUIS: Question-guided Insights Generation for Automated Exploratory Data Analysis

### v1

@@ Abstract
Abhijit Manatkar 1, Ashlesha Akella1,2, Parthivi Gupta1, Krishnasuri Narayanam1, 1IBM Research, India, 2Indian Institute of Technology Kharagpur, India,
@@ 1 Introduction
ADE using reinforcement learning is studied Milo and Somech (2018a); Bar El et al. (2019, 2020); Personnaz et al. (2021); Garg et al. (2023) to achieve full automation. While these systems minimize user involvement, they often demand dataset-specific training and substantial computational resources, particularly as the number of features, categorical values, or patterns increases, making the process increasingly challenging.
@@ 2 Preliminaries
The QUIS system incorporates the following insight types as candidates for our patterns: (1) Trend - An increasing or decreasing trend is seen in a set of values. (2) Outstanding Value - The largest (or smallest) value in a set of values is significantly larger (or smaller) than all other values in the set. (3) Attribution - The highest value accounts for a large proportion ($\geq 50\%$) of the total of all values in the set. (4) Distribution Difference - The distribution of values in a set changes notably from one subspace to another.
@@ 4 Experimental Evaluation
In our study, we evaluated the QUIS pipeline’s effectiveness using human assessment and insight scores on three datasets: Sales Verma (2024), Adidas Sale Chaudhari (2022) and Employee Attrition Subhash (2017). Human evaluation focused on the individual insights assessing Relevance, Comprehensibility, and Informativeness (details in Appendix C). We tested two conditions: 1. OnlyStats, replacing the QUGen module with a purely statistics based card generation module, to assess the autonomous performance of ISGen, 2. QUIS, where both QUGen and ISGen were involved. For further information about the parameters of the experimental conditions, please refer to Appendix D. The insights were evaluated by six participants who are well-versed in data analysis, with each insight assessed by three different evaluators. Each criterion - relevance, comprehensibility, and informativeness - was rated on a scale of 1 to 5; where 1 indicated the insight was not relevant, comprehensible, or informative; and 5 indicated the insight was highly relevant, comprehensible, or informative.
@@ Appendix A Scoring Functions for ISGen
    As this pattern holds when the highest value is more than 50% of the total, the threshld is set as $T_{\texttt{Attr}}=0.5$.

### v3

@@ Abstract
Abhijit Manatkar IBM Research, India abhijitmanatkar@ibm.com                        Ashlesha Akella IBM Research, India ashlesha.akella@ibm.com
Parthivi Gupta thanks: Work done as part of internship at IBM Research, India. Indian Institute of Technology Kharagpur parthivig@kgpian.iitkgp.ac.in                        Krishnasuri Narayanam IBM Research, India knaraya3@in.ibm.com
@@ 1 Introduction
ADE using reinforcement learning is studied Milo and Somech (2018a); Bar El et al. (2019, 2020); Personnaz et al. (2021); Garg et al. (2023); Manatkar et al. (2024) to achieve full automation. While these systems minimize user involvement, they often demand dataset-specific training and substantial computational resources, particularly as the number of features, categorical values, or patterns increases, making the process increasingly challenging.
@@ 2 Preliminaries
The QUIS system incorporates the following insight types as candidates for our patterns:
  1. 1.
    Trend - An increasing or decreasing trend is seen in a set of values.
  2. 2.
    Outstanding Value - The largest (or smallest) value in a set of values is significantly larger (or smaller) than all other values in the set.
  3. 3.
    Attribution - The highest value accounts for a large proportion ($\geq 50\%$) of the total of all values in the set.
  4. 4.
    Distribution Difference - The distribution of values in a set changes notably from one subspace to another.
@@ 3.1 Question Generation (QUGen)
QUGen prompts the language model in a structured way to generate the Breakdown and Measure components, conditioning them on the Reason and Question. This follows the Chain-of-thought prompting approach Wei et al. (2022), where the Reason and Question express the analysis intent behind each insight, ensuring the insights have stronger semantic justification and coherence.
@@ 4 Experimental Evaluation
In our study, we evaluated the QUIS pipeline’s effectiveness using human assessment and insight scores on three datasets: Sales Verma (2024), Adidas Sale Chaudhari (2022) and Employee Attrition Subhash (2017). Human evaluation focused on the individual insights assessing Relevance, Comprehensibility, and Informativeness (details in Appendix C). We tested two conditions:
  1. 1.
    OnlyStats, replacing the QUGen module with a purely statistics based card generation module, to assess the autonomous performance of ISGen
  2. 2.
    QUIS, where both QUGen and ISGen were involved.
Replicating prior work to establish robust baselines Ma et al. (2023); Guo et al. (2024); Weng et al. (2024) is challenging due to the lack of available code, datasets, and implementation details. Additionally, the differences in insight types and presentation formats across existing approaches make direct comparisons difficult. Therefore, our main focus is on comparing QUIS, against the baseline OnlyStats. For further information about the parameters of the experimental conditions, please refer to Appendix D.
The insights were evaluated by six participants who are well-versed in data analysis, with each insight assessed by three different evaluators. Each criterion - relevance, comprehensibility, and informativeness - was rated on a scale of 1 to 5; where 1 indicated the insight was not relevant, comprehensible, or informative; and 5 indicated the insight was highly relevant, comprehensible, or informative.
@@ Appendix A Scoring Functions for ISGen
    As this pattern holds when the highest value is more than 50% of the total, the threshold is set as $T_{\texttt{Attr}}=0.5$.